OpenStudy (anonymous):

the patel manufacturing company allows for a defect rate of 0.6% on its production line. the defect average for last week was 4 out of 4,400. what is the percent difference between the allowed defeat rate and the average defeat rate for last week?

OpenStudy (dumbcow):

4/4400 = 1/1100 = 0.00090909 = .090909% difference = .6 - .090909